# HandOS: 3D Hand Reconstruction in One Stage

### v1

@@ 2 Related Work
Figure 2: Overview of HandOS framework. Left: overall architecture. Right: interactive 2D-3D decoder. With off-the-shelf features, bounding boxes, and category scores from a frozen detector, the interactive 2D-3D decoder, including query filtering, expansion, lifting, and interactive layers, can understand hand pose and shape via estimating keypoints in both 2D and 3D spaces.
@@ 3.3 Interactive 2D-3D Decoder
The reference boxes for 2D joints can also be derived from instances following EDPose [66]:
@@
Additionally, a third query transformation is employed, namely query lifting, wherein queries and reference boxes are elevated from 2D joints to 3D vertices and camera translation. To this end, we design a lifting matrix $\mathbf{L}\in\mathbb{R}^{(V+1)\times J}$ as the weights for the linear combination between 2D and 3D queries, which is initialized with MANO skinning weights. Based on $\mathbf{L}$, the lifting process can be formulated as
@@
$[\mathbf{B}^{\mathbf{V}},\mathbf{B}^{\mathbf{t}}]\in\mathbb{R}^{K\times(V+1)\times 4}=\mathbf{L}\diamond\mathbf{B}^{\mathbf{J}^{2D}},$
$[\mathbf{B}_{c}^{\mathbf{V}},\mathbf{B}_{c}^{\mathbf{t}}]=\texttt{FFN}(\mathbf{[}\mathbf{Q}^{\mathbf{V}},\mathbf{Q}^{\mathbf{t}}])+[\mathbf{B}_{c}^{\mathbf{V}},\mathbf{B}_{c}^{\mathbf{t}}],$
$[\mathbf{B}_{s}^{\mathbf{V}},\mathbf{B}_{s}^{\mathbf{t}}]=[\mathbf{B}_{s}^{\mathbf{V}},\mathbf{B}_{s}^{\mathbf{t}}]\cdot\mathbf{e}^{\mathbf{R}}_{3D},$
@@ 5 Conclusion
We leave limitations and future works in suppl. material.
Acknowledgment This work was partly supported by the National Natural Science Foundation of China under Grant 62403012 and the Postdoctoral Innovative Talent Support Program under Grant BX2023004. The author team would like to acknowledge Ling-Hao Chen for his constructive suggestions and discussions.

### v2

@@ Abstract
Abstract
This is the supplementary document of HandOS, including implementation details (Section VI), metrics (Section VII), discussion on left-right classification (Section VIII), detector adaption (Section IX), and HO3D results (Section X), as well as more comparison (Section XI), efficiency analysis (Section XII), and visual results (Section XIV). Finally, failure cases (Section XIII) and limitations are analyzed (Section XVI).
@@ 2 Related Work
Figure 2: Overview of HandOS framework. Left: overall architecture. Right: interactive 2D-3D decoder. With off-the-shelf features, bounding boxes, and category scores from a frozen detector, the interactive 2D-3D decoder, including query filtering, expansion, lifting, and interactive layers, can understand hand pose and shape via estimating keypoints in both 2D and 3D spaces. Each query $\mathbf{Q}$ is associated with a reference box, which is not depicted in the figure for conciseness.
@@ 3.3 Interactive 2D-3D Decoder
The reference boxes for 2D joints $\mathbf{R}^{\mathbf{J}^{2D}}\in\mathbb{R}^{K\times J\times 4}$ can also be derived from instances following EDPose [66]:
@@
Additionally, a third query transformation is employed, namely query lifting, wherein queries and reference boxes are elevated from 2D joints to 3D vertices and camera translation. To this end, we design a learnable lifting matrix $\mathbf{L}\in\mathbb{R}^{(V+1)\times J}$ as the weights for the linear combination between 2D and 3D queries, which is initialized with MANO skinning weights. Based on $\mathbf{L}$, the lifting process can be formulated as
@@
$[\mathbf{R}^{\mathbf{V}},\mathbf{R}^{\mathbf{t}}]\in\mathbb{R}^{K\times(V+1)\times 4}=\mathbf{L}\diamond\mathbf{R}^{\mathbf{J}^{2D}},$
$[\mathbf{R}_{c}^{\mathbf{V}},\mathbf{R}_{c}^{\mathbf{t}}]=\texttt{FFN}(\mathbf{[}\mathbf{Q}^{\mathbf{V}},\mathbf{Q}^{\mathbf{t}}])+[\mathbf{R}_{c}^{\mathbf{V}},\mathbf{R}_{c}^{\mathbf{t}}],$
$[\mathbf{R}_{s}^{\mathbf{V}},\mathbf{R}_{s}^{\mathbf{t}}]=[\mathbf{R}_{s}^{\mathbf{V}},\mathbf{R}_{s}^{\mathbf{t}}]\cdot\mathbf{e}^{\mathbf{R}}_{3D},$
@@ 5 Conclusion
Acknowledgment This work was partly supported by the National Natural Science Foundation of China under Grant 62403012, Grant 62233001, Grant U23B2037 and the Postdoctoral Innovative Talent Support Program under Grant BX2023004. The authors acknowledge Ling-Hao Chen for constructive discussions.